OpenStudy (gamenerd123):

Solve for a. 2(a+4)+6a=48

OpenStudy (gamenerd123):

a = ?

likeabossssssss (likeabossssssss):

is there answer?

OpenStudy (skullpatrol):

http://www.ck12.org/student/ Please look at^

OpenStudy (gamenerd123):

How do I use this site?

RhondaSommer (rhondasommer):

distribute the 2\[2(a+4)+6a=48\\2a+8+6a=48\]add 6a and 2a\[2a+8+6a=48\\8a+8=48\]subract 8 on both sides of the equation.\[8a+8=48\\8a+8-8=48=8\\8a=40\]divide 8 on both side of the equation\[8a=40\\\frac{8x}{8}=\frac{40}{8}\\x=5\]

OpenStudy (queenbb):

hi i need some help on my question plz

OpenStudy (gamenerd123):

Ahh okay

OpenStudy (gamenerd123):

Just one more question and I'll be with you Queen